# Nonparametric Jackknife Instrumental Variable Estimation and Confounding Robust Surrogate Indices

### v1

@@ 3.6 Point identification and strong identification
We say that $\theta(h^{\star})$ is point identified if for any solution $h$ of the CMR, that is for any $h\in\mathcal{H}_{0}$, $\theta(h)=\theta(h^{\star})$.
Let $\alpha$ be the importance sampling ratio between the distribution of $S$ in the target treatment arm ($A=a^{\prime}$) and in the historical experiments ($A\in[K]$), that is
$\alpha(s)=\frac{p_{S(K+1)}(s)}{\frac{1}{K}\sum_{a=1}^{K}p_{S(a)}(s)}=\frac{p_{S^{(K+1)}}(s)}{p_{S}(s)}.$ (7)
For any $h$ we can rewrite $\theta(h)$ as an expectation under the historical data experiment in terms of the importance sampling ratio:
$\theta(h)=\mathbb{E}[\alpha(S)h(S)].$ (8)
The function $\alpha$ is then the Riesz representer of the functional $\theta$. It is then immediate that $\theta(h^{\star})$ is point identified if, and only if, $\alpha\in\mathcal{N}(\Pi)^{\perp}=\overline{\mathcal{R}(\Pi^{*})}$.
@@
We say that $\theta(h^{\star})$ is strongly identified if $\alpha\in\mathcal{R}(\Pi^{*}\Pi)$.
Note that strong identification trivially implies identification. Let $\Xi_{0}$ be the set of functions $\xi:\mathcal{S}\to\mathbb{R}$ such that $\alpha=\Pi^{*}\Pi\xi$. Strong identification is equivalent to $\Xi_{0}\neq\emptyset$. Under strong identification, $\theta(h^{\star})$ admits an additional representation, as we make explicit in the next lemma.
@@
$\left\lvert\mathbb{E}[\psi(h,\xi)]-\theta^{\star}\right\rvert\leq\left\lVert\Pi(\xi-\xi_{0})\right\rVert\times\left\lVert\Pi(h-h_{0})\right\rVert,$ (10)
@@
$\overline{\mathcal{R}(\Pi^{*})}=\mathcal{R}(\Pi^{*})=\mathcal{R}(\Pi^{*}\Pi).$ (11)
4 Estimation of the debiasing nuisance function
Risk for debiasing nuisance.
Observe that $\xi\in\Xi_{0}$, that is $\Pi^{*}\Pi\xi=\alpha$, is the first order condition in the minimization w.r.t. functions $L_{2}(\mathcal{S})$ of the risk
$R_{2}(\xi)=$ $\frac{1}{2}\mathbb{E}\left[[\Pi\xi](A)^{2}\right]-\mathbb{E}\left[\alpha(S)\xi(S)\right]$ (12)
@@ 4 Estimation of the debiasing nuisance function
$=$ $\frac{1}{2}\mathbb{E}\left[[\Pi\xi](A)^{2}\right]-\mathbb{E}\left[\xi(S^{(K+1)})\right].$ (13)
Similarly to the case of the primary nuisance function $h$, we introduce a penalized cross-fold empirical analog of the population debiasing risk. Specifically, we let
$\widehat{R}_{2,\mu}(\xi)=\frac{1}{2K}\sum_{a=1}^{K}[\widehat{\Pi}_{0}\xi](a)[\widehat{\Pi}_{1}\xi](a)-\frac{1}{n^{\prime}}\sum_{A_{i}=a^{\prime}}\xi(S^{(K+1)}_{i})+\mu\|\xi\|_{2,N}^{2}$ (14)
Let $\widehat{\xi}$ be a minimizer of $\widehat{R}_{2,\mu}(\xi)$, and let $\xi_{0}$ be the minimum norm element of $\Xi_{0}\cap\mathcal{H}$. Our convergence rate results require the following condition.
Assumption 5 (Realizability for $\xi$).
It holds that $\Xi_{0}\cap\mathcal{H}\neq 0$.
Slow rate.
Let $\mathcal{R}^{\prime}_{n^{\prime}}(\mathcal{H},\delta)=E[\sup_{\xi\in\mathcal{H}:\|\xi\|_{(K+1)}\leq\delta}\frac{1}{n^{\prime}}\sum_{i=1}^{n^{\prime}}\epsilon_{i}\xi(S^{\prime}_{i})]$ be the localized Rademacher complexity of $\mathcal{H}$ on the new dataset ($A=K+1$). Let $(r^{\prime}_{n^{\prime}})^{2}$ be an upper bound on $\mathcal{R}^{\prime}_{n^{\prime}}(\mathcal{H},1)$.
@@
Without any further assumptions, we obtain the following slow rate result.
Theorem 5 (Slow rate for $\xi$).
@@
Let $\zeta>0$. Suppose that assumption 5 holds. Suppose that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$, and $\sqrt{\log\log\delta_{N}/N}=o(\delta_{N})$. Then it holds with probability $1-O(\zeta)$ that
$\|\Pi(\xi-\xi_{0})\|=\widetilde{O}(\delta_{N}^{1/2}+r^{\prime}_{n^{\prime}}).$ (15)
Fast rate.
Let $\delta^{\prime}_{n^{\prime}}$ be a solution to $\mathcal{R}^{\prime}_{n^{\prime}}(\mathcal{H},\delta^{\prime})\leq(\delta^{\prime}_{n^{\prime}})^{2}$. We can obtain a fast rate under the following so-called source condition on the Riesz representer $\alpha$.
@@
We say that the source condition with parameter $\beta$ holds for $\alpha$ if $\alpha=(\Pi^{*}\Pi)^{\beta}w_{0}$ for some $w_{0}:\mathcal{S}\to\mathbb{R}$.
@@
Let $\zeta>0$.Suppose that assumption 5 holds. Suppose that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$, and $\sqrt{\log\log\delta_{N}/N}=o(\delta_{N})$. Further suppose that the source condition holds with parameter $\beta\in(1,2]$ and that $w_{0}$ in the source condition is such that $\sup_{K}\|w_{0}\|<\infty$. Set $\lambda=\delta_{n^{\prime}}^{\frac{2}{2+\beta}}$. Then it holds with probability $1-O(\zeta)$ that
$\|\Pi(\widehat{\xi}-\xi_{0})\|=\widetilde{O}((\delta^{\prime}_{n^{\prime}})^{-\frac{1+\beta}{2+\beta}}).$ (16)
@@
$\|\xi-\xi_{0}\|=\widetilde{O}((\delta^{\prime}_{n^{\prime}})^{-\frac{\beta}{2+\beta}}).$ (17)
Note that the convergence rates above are in terms of the norm induced by historical data distribution.

### v2

@@ 3.6 Identification and strong identification
Let $r_{0,K}$ be the $[K]\to\mathbb{R}$ function defined for every $a\in[K]$ by $r_{0,K}(a)=E[Y\mid A=a]$. We say that $\theta(h^{\star})$ is identified if for any $h\in\mathcal{H}$ such that $\mathcal{T}_{K}h=r_{0,K}$, that is for any $h\in\mathcal{H}_{0,K}\cap\mathcal{H}$, $\theta(h)=\theta(h^{\star})$.
Let $\rho$ be the importance sampling ratio between the distribution of $S$ in the target treatment arm ($A=0$) and in the historical experiments ($A\in[K]$), that is, for every $s\in\mathcal{S}$,
$\rho_{K}(s)=\frac{p_{S(\mathrm{new})}(s)}{\frac{1}{K}\sum_{a=1}^{K}p_{S(a)}(s)}=\frac{p_{S^{\mathrm{new}}}(s)}{p_{S}(s)}.$ (6)
For any $h\in\mathcal{H}$ we can rewrite $\theta(h)$ as an expectation under the historical experiments data in terms of the importance sampling ratio:
$\theta(h)=\mathbb{E}[\rho_{K}(S)h(S)]=\mathbb{E}[\alpha_{K}(S)h(S)]\qquad\text{where}\qquad\alpha_{K}=\Pi(\rho_{K}\mid\mathcal{H}),$ (7)
where $\Pi(\cdot\mid\mathcal{H})$ is the $L_{2}$ projection on $\mathcal{H}$. The function $\alpha_{K}$ is then the Riesz representer of the functional $\theta$. It is then immediate that $\theta(h^{\star})$ is point identified if, and only if, $\alpha_{K}\in\mathcal{N}(\mathcal{T}_{K})^{\perp}=\overline{\mathcal{R}(\mathcal{T}_{K}^{*})}$. Note that identification in our setting is a strong condition: indeed, $\mathcal{N}(\mathcal{T}_{K})^{\perp}$ is a $K$-dimensional subspace of $\mathcal{H}$, which places considerable restrictions on the allowable complexity of $\alpha_{K}$. In the asymptotic regime we consider, $K\to\infty$, which mitigates asymptotically this stringent condition. In the next subsection, we consider an approximate notion of identification which might be more realistic in our setting. Before presenting this notion, we nevertheless discuss how identification relates to another, stronger notion of identification discussed in various recent articles [ref needed].
@@
We say that $\theta(h^{\star})$ is strongly identified if $\alpha_{K}\in\mathcal{R}(\mathcal{T}_{K}^{*}\mathcal{T}_{K})$, that is if $\Xi_{K}:=\{\xi\in\mathcal{H}:\mathcal{T}_{K}^{*}\mathcal{T}_{K}\xi=\alpha_{K}\}\neq\emptyset$.
Note that strong identification trivially implies identification. Under strong identification, $\theta(h^{\star})$ admits an additional representation, as we make explicit in the next lemma.
@@
$\left\lvert\mathbb{E}[\psi(h,\xi)]-\theta^{\star}\right\rvert\leq\left\lVert\mathcal{T}_{K}(\xi-\xi_{K})\right\rVert\times\left\lVert\mathcal{T}_{K}(h-h_{K})\right\rVert,$ (9)
@@
We refer to $\xi_{K}$ above as the debiasing nuisance and to $h_{K}$ as the primary nuisance. As we will see in section 5, the mixed bias property implies that we can get parametric-rate inference for $\theta^{\star}$ from nonparametrically estimated nuisances as long as the product of their weak norms converges to zero at a parameteric rate.
@@
$\overline{\mathcal{R}(\mathcal{T}_{K}^{*})}=\mathcal{R}(\mathcal{T}_{K}^{*})=\mathcal{R}(\mathcal{T}_{K}^{*}\mathcal{T}_{K}).$ (10)
Thefore, if $\theta^{*}$ is identified, the existence of a debiasing nuisance and the double robust representation is automatically guaranteed.
3.7 Approximate strong identification
In this subsection, we provide a notion of approximate identification which accounts for violations of the stringent identification condition of the previous section.
Definition 3.
We say that $\theta(h^{\star})$ is $(K,\delta)$-approximately strongly identified if $\min_{q\in\mathcal{Q}_{K}}\|\alpha_{K}-\mathcal{T}_{K}^{*}q\|\leq\delta$.
Remark 3.
Using the same arguments as in the proof of theorem 4, we can show that $(K,\delta_{K})$-strong identification is equivalent to $\inf_{\xi\in\mathcal{H}}\|\alpha_{K}-\mathcal{T}_{K}^{*}\mathcal{T}_{K}\xi\|\leq\delta_{K}$, which makes more explicit the connection with strong identification.
Remark 4.
As $\mathcal{R}(\mathcal{T}^{*}_{K})$ grows with $K$, the sequence of $(\delta_{K})$ of smallest values of $\delta$ for which approximate strong identification holds is non-increasing. Intuitively, it should decrease faster if the historical experiments impose more “complementary” moment restrictions.
The following theorem shows that under approximate strong identification, $\theta(h^{\star})$ is identified up to an error that decreases with $K$.
Theorem 5 (Approximate target identification).
If $\|\Pi(h^{\star}\mid\mathcal{N}(\mathcal{T}_{K}))\|\leq\epsilon_{K}$, $(K,\delta_{K})$-approximate strong identification holds and $q_{K}\in\operatornamewithlimits{argmin}\|\alpha_{K}-\mathcal{T}_{K}^{*}q\|$ and $h_{K}^{\dagger}$ is the minimum norm element of $\mathcal{H}_{0,K}\cap\mathcal{H}$, then, for $\psi(h,q)=h(S^{\mathrm{new}})+q(A)(Y-h(S))$, we have
$|\mathbb{E}[\psi(h_{K}^{\dagger},q_{K})]-\theta(h^{\star})|\leq\epsilon_{K}\delta_{K}.$ (11)
We obtain the following appproximate version of the mixed-bias result as a corollary.
Corollary 1.
Let $h\in\mathcal{H}$, $q\in\mathcal{Q}$. Let $q_{K}^{\dagger}$ be the minimum norm element of $\operatornamewithlimits{argmin}_{q\in\mathcal{Q}}\|\alpha_{K}-\mathcal{T}_{K}^{*}q\|$. Under the assumptions and notation of the previous theorem,
$|\mathbb{E}\psi(h,q)-\theta(h^{\star})|\leq\epsilon_{K}\delta_{K}+\delta_{K}\|h-h_{K}^{\dagger}\|+\|q-q_{K}^{\dagger}\|\|\mathcal{T}_{K}(h-h^{\star})\|.$ (12)
Representation of $q_{K}^{\dagger}$.
While the above definition of $q_{K}^{\dagger}$ shows how to identify it from the historical data distribution and the novel data distribution, doing so requires computing several density ratios: $\alpha_{K}$ is the projection of an importance sampling ratio, and the adjoint operator $\mathcal{T}_{K}^{*}$ can be shown to involve the conditional densities of $S$ given values of $A$ in the historical data. As density ratios are typically hard to estimate this doesn’t suggest a straightforward estimation strategy. We now show that we can obtain $q_{K}^{\dagger}$ by solving two consecutive quadratic optimization problems that have easy-to-implement sample-level counterparts. We have that $\mathcal{Q}$ is the linear span of the functions $q_{a}^{*}:=\mathcal{T}_{K}^{*}\bm{1}\{\cdot=a\}$ for $a=1,\ldots,K$. Using the adjoint property, we have that
$q_{a}^{*}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\mathbb{E}[h^{2}(S)]-2E[h(S)\mid A=a].$ (13)
Finding $q_{K}^{\dagger}$ now reduces to finding
$\bm{\gamma}=\operatornamewithlimits{argmin}_{\bm{\gamma}\in\mathbb{R}^{K}}E[(\sum_{a=1}^{K}\gamma_{a}q_{a}^{*}(S))^{2}]-2E[\alpha_{K}(S)\sum_{a=1}^{K}\gamma_{q}q_{a}^{*}(S)],$ (14)
that is, using the representer property of $\alpha_{K}$, as in automatic debiased machine learning (Chernozhukov et al., 2022),
$\bm{\gamma}=\operatornamewithlimits{argmin}_{\bm{\gamma}\in\mathbb{R}^{K}}\sum_{1\leq a,a^{\prime}\leq K}\gamma_{a}\gamma_{a^{\prime}}\mathbb{E}[q_{a}^{*}(S)q_{a^{\prime}}^{*}(S)]-2\sum_{a=1}^{K}\gamma_{a}E[q_{a}^{*}(S^{\mathrm{new}})].$ (15)
We discuss how to translate the above at the sample level for estimation in subsection 4.2 of the next section.
4 Estimation of debiasing nuisances
Estimation of the primary nuisance $h$ is a direct application of the npJIVE results of section 2. In this section we show how to estimate the debiasing nuisances under both identification and approximate identification. Under identification, the strategy and the analysis closely follows that of the primary nuisance. In both cases, we provide cross-folded estimator and/or Jackknifed estimators.
4.1 Estimation of the debiasing nuisance function under (strong) identification
In this section we suppose that $\theta^{\star}$ is identified, as defined in section 3. From theorem 4, there exists a debiasing nuisance $\xi_{K}$ that is $\mathcal{T}_{K}^{*}\mathcal{T}_{K}\xi_{K}=\alpha_{K}$. Observe that $\mathcal{T}_{K}^{*}\mathcal{T}_{K}\xi_{K}=\alpha_{K}$ is the first order condition of the optimization problem $\min_{\xi\in\mathcal{H}}R_{2}(\xi)$ where
$R_{2}(\xi)=\mathbb{E}\left[[\mathcal{T}_{K}\xi](A)^{2}\right]-2\mathbb{E}\left[\alpha_{K}(S)\xi(S)\right].$ (16)
From the definition of $\alpha_{K}$, for any $\xi\in\mathcal{H}$, we can rewrite $R_{2}(\xi)$ as
$R_{2}(\xi)=\mathbb{E}\left[[\mathcal{T}_{K}\xi](A)^{2}\right]-2\mathbb{E}\left[\xi(S^{\mathrm{new}})\right],$ (17)
which avoids reference to typically hard to estimate density ratios, as in automatic debiased maching learning (Chernozhukov et al., 2022).
As in the case of the primary nuisance, the risk for the debiasing nuisance involves the quadratic projected term $\mathbb{E}\left[[\mathcal{T}_{K}\xi](A)^{2}\right]$, which, as we argued in section 2 cannot be estimated consistently by $K^{-1}\sum_{a=1}^{K}[\widehat{\mathcal{T}}_{K}\xi](a)$ if $n$ stays bounded. To overcome this issue, as in section 2, we introduce a Tikhonov regularized cross-fold empirical analog of the population debiasing risk. Specifically, for any regularization level $\mu>0$, we let
$\widehat{R}_{2,\mu}(\xi)=\frac{1}{2K}\sum_{a=1}^{K}[\widehat{\mathcal{T}}_{K,0}\xi](a)[\widehat{\mathcal{T}}_{K,1}\xi](a)-\frac{1}{n_{\mathrm{new}}}\sum_{A_{i}=a^{\prime}}\xi(S^{\mathrm{new}}_{i})+\mu\|\xi\|_{2,N}^{2}.$ (18)
Let $\widehat{\xi}$ be a minimizer over $\mathcal{H}$ of the regularized empirical risk $\widehat{R}_{2,\mu}(\xi)$, and let $\xi_{K}^{\dagger}$ be the minimum norm element of the set $\Xi_{K}$ of debiasing nuisances. We present next convergence rate guarantees for $\widehat{\xi}$. We start with a slow rate result. Before stating it, we need to introduce an additional measure of complexity: let
@@ 4.1 Estimation of the debiasing nuisance function under (strong) identification
$\mathcal{R}^{\mathrm{new}}(\mathcal{H},\delta)=\mathbb{E}\left[\sup_{\xi\in\mathcal{H}:\|\xi\|_{\mathrm{new}}\leq\delta}\frac{1}{n_{\mathrm{new}}}\sum_{i=1}^{n_{\mathrm{new}}}\epsilon_{i}\xi(S_{i}^{\mathrm{new}})\right]$ (19)
@@
be the localized Rademacher complexity of $\mathcal{H}$ on the new dataset ($A=\mathrm{new}$). Let $(r^{\mathrm{new}}_{n_{\mathrm{new}}})^{2}$ be an upper bound on $\mathcal{R}^{\mathrm{new}}(\mathcal{H},1)$. Without any further assumptions, we obtain the following slow rate result.
Theorem 6 (Slow rate for $\widehat{\xi}$).
@@
Let $\zeta>0$. Suppose that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$, and $\sqrt{\log\log\delta_{N}/N}=o(\delta_{N})$. Then it holds with probability $1-O(\zeta)$ that
$\|\mathcal{T}_{K}(\xi-\xi_{0})\|=\widetilde{O}(\delta_{N}^{1/2}+r^{\mathrm{new}}_{n_{\mathrm{new}}}+\sqrt{\mu}).$ (20)
We show next that, under an additional assumption, we can obtain a fast rate in that replaces the unlocalized Rademacher complexity $r^{\mathrm{new}}_{n_{\mathrm{new}}}$ with a critical radius of $\mathcal{H}$ over the new data set. The additional assumption we require is a so-called source condition on the Riesz representer $\alpha_{K}$, initially introduced in Bennett et al. (2023b). We state it below.
@@
We say that the source condition with parameter $\beta$ holds for $\alpha_{K}$ if $\alpha_{K}=(\mathcal{T}_{K}^{*}\mathcal{T}_{K})^{\beta}w_{K}$ for some $w_{K}\in\mathcal{H}$.
For $\beta=1$, the source condition reduces to the strong identification condition we introduced in section 3. As we will make explicit in the next theorem, the source condition buys us additional convergence speed if $\beta>1$. Let $\delta^{\mathrm{new}}_{n_{\mathrm{new}}}$ be a critical radius of $\mathcal{H}$ over the new data set defined as a solution to $\mathcal{R}^{\mathrm{new}}(\mathcal{H},\delta)\leq\delta^{2}$.
@@
Let $\zeta>0$. Suppose that $\sqrt{\log(K/\zeta)}=o(\widetilde{\delta}_{n})$, and $\sqrt{\log\log\delta_{N}/N}=o(\delta_{N})$. Further suppose that the source condition holds with parameter $\beta\in(1,2]$ and that $w_{K}$ in the source condition is uniformly bounded in $K$, that is $\sup_{K}\|w_{K}\|<\infty$. Set $\mu=(\delta^{\mathrm{new}}_{n_{\mathrm{new}}})^{\frac{2}{2+\beta}}$. Then it holds with probability $1-O(\zeta)$ that
$\|\mathcal{T}_{K}(\widehat{\xi}-\xi_{0})\|=\widetilde{O}((\delta^{\mathrm{new}}_{n_{\mathrm{new}}})^{-\frac{1+\beta}{2+\beta}}).$ (21)
@@
$\|\widehat{\xi}-\xi_{K}\|=\widetilde{O}((\delta^{\mathrm{new}}_{n_{\mathrm{new}}})^{-\frac{\beta}{2+\beta}}).$ (22)
Note that the strong norm convergence rate above is in terms of the norm induced by the historical data distribution.
4.2 Estimation of the debiasing nuisance function under approximate identification
In this section we propose an estimator of the debiasing nuisance $q_{K}^{\dagger}$ under approximate identification. While we leave guarantees to follow-up work, we heuristically justify its construction in this subsection and empirically validate its performance in section 6.
Based on the representation of $q_{K}^{\dagger}$ as the solution to two consecutive optimization problems (13) and (15), a natural estimation strategy is to define, for every $a\in[K]$,
$\widehat{q}_{a}^{*}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\frac{1}{N}\sum_{i=1}^{N}h(S_{i})^{2}-2\frac{1}{n}\sum_{i:A_{i}=a}h(S_{i}),$ (23)
and let $\widehat{q}=\sum_{a=1}^{K}\widehat{\gamma}_{a}\bm{1}(\cdot=a)$ where
$\widehat{\bm{\gamma}}=\operatornamewithlimits{argmin}_{\bm{\gamma}\in\mathbb{R}^{K}}\sum_{1\leq a,a^{\prime}\leq K}\gamma_{a}\gamma_{a^{\prime}}\frac{1}{N}\sum_{i=1}^{N}\widehat{q}_{a}^{*}(S_{i})\widehat{q}_{a^{\prime}}^{*}(S_{i})-2\sum_{a=1}^{K}\gamma_{a}\frac{1}{n_{\mathrm{new}}}\sum_{i=1}^{n_{\mathrm{new}}}\widehat{q}_{a}^{*}(S_{i}^{\mathrm{new}}).$ (24)
For the same reason we already mentioned earlier, the above risk isn’t consistent under weak-instrument asymptotics: since $\widehat{q}_{a}^{*}$ optimizes a risk with an empirical mean over $n$ terms, $\|\widehat{q}_{a}^{*}-q_{a}^{*}\|^{2}$ doesn’t shrink as $n$ stays fixed and appears in the expectation on the empirical risk in (24).
We therefore propose a Jackknifing device to eliminate squares of errors that do not converge to zero as $n$ stays fixed. For every $a\in[K]$ and $i$ in historical treatment arm $a$ (that is such that $A_{i}=a$), we define
$\widehat{q}_{a,-i}^{*}=\operatornamewithlimits{argmin}_{h\in\mathcal{H}}\frac{1}{N}\sum_{j=1}^{N}h(S_{j})-\frac{2}{n-1}\sum_{j:\begin{subarray}{c}A_{j}=a\\ j\neq i\end{subarray}}h(S_{j}).$ (25)
Now observe that for any $a$, $\mathbb{E}[q_{a}^{*}(S)^{2}]=\mathbb{E}[[\mathcal{T}_{K}q_{a}^{*}](A)q_{a}(A)]=K^{-1}\mathbb{E}[q_{a}^{*}(S)\mid A=a]$. This suggests estimating the expectations of squares in (15) with $K^{-1}n^{-1}\sum_{i:A_{i}=a}\widehat{q}_{a,-i}^{*}(S_{i})$. Concerning the cross terms $\mathbb{E}[q_{a}^{*}(S)q_{a^{\prime}}^{*}(S)]$ for $a\neq a^{\prime}$, the treatment-arm-level means that lead to $\widehat{q}_{a}$ and $\widehat{q}_{a^{\prime}}^{*}(S)$ are over different observations, so these terms shouldn’t cause issues. This discussion suggests defining $\widehat{q}=\sum_{a=1}^{K}\widehat{\gamma}_{a}\bm{1}\{\cdot=a\}$ where
$\bm{\gamma}=\sum_{1\leq a,a^{\prime}\leq K}\gamma_{a}\gamma_{a^{\prime}}C_{a,a^{\prime}}-2\sum_{a=1}^{K}\gamma_{a}\frac{1}{n_{\mathrm{new}}}\sum_{i=1}^{n_{\mathrm{new}}}\widehat{q}_{a}^{*}(S_{i}^{\mathrm{new}}),$ (26)
with
$C_{a,a^{\prime}}=\begin{cases}\frac{1}{Kn}\sum_{i:A_{i}=a}\widehat{q}_{a,-i}^{*}(S_{i})&\text{ if }a=a^{\prime},\\ \frac{1}{N}\sum_{i=1}^{N}\widehat{q}_{a}^{*}(S_{i})\widehat{q}_{a^{\prime}}^{*}(S_{i})&\text{ if }a\neq a^{\prime}.\end{cases}$ (27)